# A Mixed-Methods Study on the Implications of Unsafe Rust for Interoperation, Encapsulation, and Tooling

### v3

@@ 6.2. RQ1 - Interoperation
The majority of interview participants and 70% of survey respondents used unsafe code to call foreign functions. The five most popular languages that respondents interoperated with were C (90%), C++ (57%), JavaScript or TypeScript (22%), assembly (21%), and Python (17%). With the exception of Assembly, the relative popularity of each language is similar to results from Fulton et al. (Fulton et al., 2021). Assembly would have been included as part of the category "Other" in their survey, which was selected by only 7% of respondents. Additionally, 72% of our respondents had converted part or all of an application written in another language to Rust, and 16% of these respondents had not called foreign functions. This supports Zeng and Crichton (Zeng and Crichton, 2018)’s notion that some developers may prefer to rewrite entire applications in Rust instead of relying on interoperation to incrementally translate them. However, none of our evidence suggests that this is a dominant preference.
Memory Management
@@
Rust’s standard library provides several types that can be used to manage ownership of heap allocations. Each of these types can be converted to raw pointers when necessary, but similar to Rust’s safe references, they have distinct invariants that must be maintained to avoid undefined behavior. Participants frequently used these types to allocate memory for foreign function calls, converting them into raw pointers in the process. In particular, 54% of survey respondents who had called foreign functions also used Rust’s memory container types for this purpose. Rust’s thread-safe Arc—an atomic, reference-counted allocation—was used by 28% of these participants, while only 8% had used the non-atomic Rc. Rust’s Box provides unique ownership over a heap allocation, and it was used by 95% of these participants to allocate memory for foreign calls. One interview participant used Box extensively when porting an online multiplayer text-adventure game from C to Rust. The game’s logic was written in C, but all heap memory was allocated in Rust using Box. When a Box was unwrapped and exposed as a raw pointer to C, its address was added to a table. The table was checked and updated to ensure that each allocation was only freed once.
@@
Several interview participants mentioned having issues with Box, noting that when a Box is moved, any raw pointers to its contents are invalidated. However, participants with a background in C++ development had not expected Box to behave this way. They were accustomed to using unique_ptr from the C++ standard library, which has a similar role but different semantics:
@@
You give the raw pointer off to the interface, and then you relocate the unique_ptr, but you still rely on that pointer having valid access tags. (P0)
@@
Survey participants did not encounter this issue as frequently. Of the 51% of respondents had used Box in any unsafe context, only 17% had encountered issues when a pointer because invalid after a Box was moved. However, we did not determine if respondents had actually unwrapped a Box or otherwise retained a raw pointer to its contents, so this 17% could still represent a significant portion of the respondents who would have been able to encounter this problem.
@@
One participant observed several aliasing patterns in foreign libraries that were at odds with Rust’s restrictions. They struggled to encapsulate foreign functions that would read from a “source” pointer and write to a “destination” pointer. It was difficult to use safe references as parameters for the Rust encapsulations of these functions when the source and destination pointers needed to be aliases. The borrow checker would prevent the mutable destination reference from aliasing with the immutable source reference. Interview participants also noted that C allows unrestricted integer-to-pointer conversion, but Rust does not have an agreed-upon semantics for this behavior (Jung, 2024). Another participant observed that C libraries would often use on “a rat’s nest of pointers” (P0) to encode self-referential structures, which were difficult to represent in Rust encapsulations. Each of the aliasing violations from the most popular libraries tested by McCormack et al. (McCormack et al., 2024) were caused by incorrect encapsulations of self-referential structures.
@@
Interview participants typically used more sophisticated multithreading patterns in the Rust components of multi-language applications. One participant used granular locking within the Rust codebase, while the C codebase had a singular global lock. For another participant, the C++ component of their application was structured along a single main thread, while the Rust component used multithreading freely—“we were throwing threads around all the time” (P0). It was easy for this participant to accidentally call into C++ from the wrong Rust thread, but they saw this as less challenging to manage than it would have been to implement the entire application in C++:
@@
Table 4. How often survey respondents would avoid passing abstract data types by value to foreign functions and converting foreign raw pointers into safe references. Percentages on the left include “Always” and “Most of the time”, while those on the right include “Sometimes” and “Never”. The number of respondents in the sample is listed in parenthesis on the righthand side.
{tblr}
width=colspec=Q[c,m]X[l,m]$Q[c,m]$ ID Question Sample Distribution
[…]
[Uncaptioned image]49%24%(112)
4b How often do you intentionally avoid converting raw pointers to memory allocated by FFI calls into safe references, such as &T or &mut T?
[Uncaptioned image]23%43%(112)
[…]
@@
Participants attempted to keep interoperation with foreign code as straightforward as possible. When declaring bindings, one participant preferred using primitive types and pointers instead of abstract data types like Option and NonNull, even though Rust allows these types to be implicitly cast into raw pointers (Community, 2024). They felt that these casts hid useful contextual information, so they preferred to use raw pointers explicitly. A few participants indicated that they sometimes avoided passing structs by value across foreign boundaries. One of these participants observed that certain Rust programs would behave incorrectly when structs were moved across a foreign boundary instead of copied, or if they had a destructor on the C++ side of the boundary. They typically avoided passing structs by value in either situation, and they recommended that “…if you’re doing FFI, the only things you should pass by value are Copy types” (P0). Other interview participants also indicated that they felt comfortable passing structs by value as long as they had copy semantics and equivalent types on each side of the boundary.
@@
We investigated these information-hiding practices using the survey questions shown in Table 0, where respondents indicated if they “intentionally avoid” passing structs by value or casting foreign pointers into Rust’s references. We added “Unsure” as a response option for these questions since they indirectly rely on participants having a notion of why they would “intentionally” avoid these behaviors. Neither practice was consistently followed, but more respondents would avoid passing abstract data types by value most or all of the time (49%, 6.2), than would avoid casting foreign pointers into references (43%, 6.2). Key Findings - RQ1 (Interoperation).
How do Rust developers reason about memory safety across foreign function boundaries? The majority of participants called foreign functions, which were typically written in C and C++. Certain foreign libraries used aliasing and concurrency patterns that conflicted with the expectations of Rust’s type system, which made them difficult to encapsulate. To make interoperation easier, participants minimized their interaction with foreign libraries, but they frequently used Rust’s memory containers as allocators, and few actively avoided casting foreign pointers into safe references. Developers identified safety properties and design patterns of foreign APIs that match the characteristics of libraries where McCormack et al. (McCormack et al., 2024) detected aliasing violations, indicating that foreign function calls may be a significant unchecked source of undefined behavior in the Rust ecosystem. We expect that developers who write encapsulations for foreign APIs would benefit from analysis tools that can infer their safety properties. This information could be leveraged to automatically generate sound, safe encapsulations.
@@ 6.3. RQ2 - Tooling
Table 5. How often survey respondents audited their dependencies, used a debugger, wrote tests, and executed tests in Miri (if they had used Miri at least once). Percentages on the left include “Always” and “Most of the time”, while those on the right include “Sometimes” and “Never”. The number of respondents in each sample is included in parenthesis on the righthand side.
{tblr}
width=colspec=Q[c,m]X[l,m]$Q[c,m]$
@@
[…]
[Uncaptioned image]61%30%(160)
Used Miri
[Uncaptioned image]72%19%(98)
5b How often do you run test cases in Miri? Used Miri
[Uncaptioned image]38%56%(98)
[…]
[Uncaptioned image]6%88%(160)
Used Auditing Tools
[Uncaptioned image]10%80%(81)
[…]
Table 6. Whether or not survey respondents who had used foreign functions would trust the correctness of foreign function bindings written by hand and generated using a tool. Percentages on the left include “Definitely yes” and “Probably yes”, while those on the right include “Probably not” and “Definitely not”. The number of respondents the each sample is included in parenthesis on the righthand side.
{tblr}
width=colspec=Q[c,m]X[l,m]$Q[c,m]$ ID Question Sample Distribution
[…]
[Uncaptioned image]45%7%(112)
6b Do you trust FFI bindings that are generated by a tool?
[Uncaptioned image]79%2%(112)
[…]
Participants used a wide variety of development tools to assist with validating their design choices. We focused primarily on dynamic tools for bug finding and static tools for generating bindings to foreign functions. However, certain participants also described their experiences using debuggers, verifiers, and auditing tools.
@@
Most participants used dynamic bug finding tools. Interview participants often mentioned Miri, AddressSanitizer, and Valgrind. We prompted survey respondents to indicate which dynamic tools they had used from a list of options mentioned by our interview participants. Survey respondents most frequently used Miri (61%), Valgrind (44%), AddressSanitizer (26%), and cargo-fuzz (26%). Each of the remaining tools listed in our survey was used by less than 15% of respondents. Three interview participants and 31% of survey respondents used some form of fuzzing tools. One interview participant leveraged industry sponsorship to fuzz a JIT compiler for hours-on-end. Another participant implemented a BTree with optimizations for high performance computing, and they used libfuzzer to perform differential testing (McKeeman, 1998) against the implementation provided by Rust’s standard library. Of the 26 developers interviewed by Höltervennhoff et al. (Höltervennhoff et al., 2023) only seven had used Miri, seven used fuzzing tools, and three used Valgrind. Our results indicate that this is not representative of a broader trend in the Rust community.
The majority of survey respondents also wrote tests for applications that used unsafe code. As shown in Table 0, 61% (6.3) of all participants wrote tests most of the time or always. This was equally true for the subset of participants who had used Miri; 72% would write test cases at least most of the time. However, 56% (6.3) of survey respondents who had used Miri only sometimes used it to run their test cases, if at all. Höltervennhoff et al. (Höltervennhoff et al., 2023) also had similar findings. The majority of their participants wrote unit tests for unsafe code, but a subset claimed that they would infrequently run these tests because foreign function calls or low-level hardware interaction made testing “impossible.” While notable, we cannot make any causal claims between Miri’s limitations and this lack of testing in our survey population. Additionally, we did not ask if survey respondents regularly ran their test cases. Interview participants frequently mentioned that Miri’s slow performance and lack of support for foreign function calls prevented them from using it to find bugs in certain categories of applications. Survey respondents who had used Miri encountered similar issues; 62% were also deterred from using it due to its lack of support for foreign functions (43%), slow performance (26%), or lack of support for inline assembly (19%). Developers interviewed by Höltervennhoff et al. (Höltervennhoff et al., 2023) also struggled with Miri’s lack of support for key features, but none had mentioned its performance.
@@
Interview participants encountered several issues with tools that generate bindings to foreign functions. Participants reported that bindgen added significant runtime overhead during build times and was not as expressive as they desired. One participant observed that bindgendid some weird stuff for anonymous unions and structs” (P0) but that this was easy to fix by modifying the tool’s configuration. However, another participant found it necessary to switch from using bindgen to writing bindings by hand, because it was not capable of expanding the macros that they had used to define core data structures. Others felt that the bindings they wrote by hand were more idiomatic than the output of a tool. The majority (58%) of survey respondents who created bindings used multiple methods, while 28% only generated bindings and 14% only wrote them manually.
Interview participants who wrote bindings by hand viewed this process as error-prone, noting that bindings become a “fiddly mess” (P0) when you need to ensure that interfaces remain consistent on each side of the foreign boundary. These errors clearly do occur in practice, as 35% of survey respondents who had called foreign function had also encountered incorrect bindings. McCormack et al. (McCormack et al., 2024) also found several instances of incorrect bindings in Rust libraries. Some were intentional—developers neglected to add return types for values they did not use—but others introduced aliasing violations or uninitialized reads. Survey respondents who had called foreign functions were somewhat more likely to trust generated bindings. As shown in Table 0, only 45% (6.3) of respondents would at least “probably” trust handwritten bindings, compared to 79% (6.3) who would trust generated bindings. However, developers were still more likely to trust handwritten bindings than to distrust them.
@@
Few interview participants reported using a debugger with Rust code. Only 10% of survey respondents used a debugger daily, 29% used one monthly, and 44% used one yearly at most, if at all. One interview participant had used both MSVC’s debugger and LLDB, and while both worked, neither matched the quality of the other parts of the Rust toolchain, “where you get all the bells and whistles that you want, and things just work nicely together” (P0). Mozilla’s RR debugger (O’Callahan et al., 2017) was useful for one participant to resolve bugs found from fuzzing, but it was generally more common for participants to report using “printf-style” (P0) debugging.
@@
Few interview participants and only 10 survey respondents used tools that apply formal methods for static and dynamic verification. Interview participants mentioned using Kani, Prusti, and Crucible. Kani was useful for one participant, who needed to prove that a particular routine would never execute more than twice for any input. Another participant who had “tried a bunch of the tools” in this category found that Prusti was the most complete, but they felt that it would be more effective if these tools were “inherently parts of Rust” so that they could “hoist a lot of currently unsafe code into safe code” (P0). Of the ten survey respondents who had used “formal methods tools,” five had used Kani, three had used Prusti, three had used Creusot, one had used Flux, and one had used Crucible. We provided a short-response option for this question to include any tools that were missing from our list.
@@
Survey respondents rarely audited their dependencies’ use of unsafe code. More than 80% (6.3) of respondents only sometimes audited their dependencies’ use of unsafe code, if at all. It is possible that a subset of the participants who answered “Never” to this question did not have any dependencies to audit. However, 51% of our respondents had used automated auditing tools, compared to four interviewed by Höltervennhoff et al. (Höltervennhoff et al., 2023). Of this subset, 63% had used cargo-audit, 42% had used cargo-update, 31% had used cargo-deny, 22% had used cargo-geiger, and only 7% had used cargo-vet.
@@
What tools do Rust developers use when contributing to applications that include unsafe code, and how could tooling be improved? Miri appears to be the “de facto” (Gäher et al., 2024) bug finding tool for the Rust ecosystem, since it was used by 61% of all respondents. However, 62% of the respondents who had used Miri before were deterred from using it again due to its performance and lack of support for key features—predominately, foreign function calls. Tools for generating foreign function bindings were often slow and required additional configuration, so a subset of developers still preferred writing bindings by hand. Participants infrequently used debuggers, and only 10 survey respondents had used formal methods tools. Survey respondents infrequently audited their dependencies, but 51% had used one or more auditing tools. Improvements to Miri may be the most effective way to provide stronger safety guarantees for the majority of our audience, since it was the most widely used tool in our survey.

### v4

@@ 5.2. RQ1 - Interoperation
The majority of interview participants and 70% of survey respondents used unsafe code to call foreign functions. The five most popular languages that respondents interoperated with were C (90%), C++ (57%), JavaScript or TypeScript (22%), Assembly (21%), and Python (17%). With the exception of Assembly, the relative popularity of each language is similar to results from Fulton et al. (Fulton et al., 2021). Their survey included Assembly as part of the "Other" category, which only 7% of their respondents had selected. Although 72% of our respondents had converted part or all of an application written in another language to Rust, 25% of these respondents had not called foreign functions. This subset might have ported these applications wholesale instead of opting for an incremental approach.
Aliasing & Memory Management
@@
Rust’s standard library provides several types that can be used to manage ownership of heap allocations. Each of these types can be converted to raw pointers when necessary, but similar to Rust’s safe references, they have distinct invariants that must be maintained to avoid undefined behavior. Participants frequently used these types to allocate memory for foreign function calls, converting them into raw pointers in the process. In particular, 54% of survey respondents who had called foreign functions also used one of Rust’s memory container types for this purpose. Rust’s thread-safe Arc—an atomic, reference-counted allocation—was used by 28% of these participants, while only 8% had used the non-atomic Rc. Rust’s Box provides unique ownership over a heap allocation, and it was used by 95% of these participants to allocate memory for foreign calls. One interview participant used Box extensively when porting an online multiplayer text-adventure game from C to Rust. The game’s logic was written in C, but all heap memory was allocated in Rust using Box. When a Box was unwrapped and exposed as a raw pointer to C, its address was added to a table. The table was checked and updated to prevent double-free errors.
@@
Several interview participants mentioned having issues with this type, noting that when a Box is moved, any raw pointers to its contents become invalid. Participants with a background in C++ development had not expected Box to behave this way. They were accustomed to using unique_ptr from the C++ standard library. This has a similar role, but different semantics:
@@
You give the raw pointer off to the interface, and then you relocate the unique pointer, but you still rely on that pointer having valid access tags. (P9)
@@
Survey participants did not encounter this issue as frequently. Of the 51% of respondents had used Box in any unsafe context, only 17% had encountered issues when a pointer became invalid after a Box was moved. However, we did not determine if respondents had actually unwrapped a Box or otherwise retained a raw pointer to its contents, so this 17% could still represent a significant portion of the respondents who would have been able to encounter this problem. Our population is most representative of experienced Rust developers, so it is possible that the average Rust developer is even more likely to have misused Rust’s Box type in these contexts.
@@
One participant observed several aliasing patterns in foreign libraries that were at odds with Rust’s restrictions. They struggled to encapsulate foreign functions that would read from a “source” pointer and write to a “destination” pointer. It was difficult to use safe references as parameters for the Rust encapsulations of these functions when the source and destination pointers needed to be aliases. The borrow checker would prevent the mutable destination reference from aliasing with the immutable source reference. Interview participants also noted that C allows unrestricted integer-to-pointer conversion, but Rust does not have an agreed-upon semantics for this behavior (Jung, 2024). Another participant observed that C libraries would often use on “a rat’s nest of pointers” (P14) to encode self-referential structures, which were difficult to represent in Rust encapsulations.
@@
This pattern matches two of the aliasing violations observed by McCormack et al. (McCormack et al., 2025b), where "unique" mutable references were cast into raw pointers and stored into the heap by foreign functions.
@@
Interview participants typically used more sophisticated multithreading patterns in the Rust components of multilanguage applications. One participant used granular locking within the Rust codebase, while the C codebase had a singular global lock. For another participant, the C++ component of their application was structured along a single main thread, while the Rust component used multithreading freely—“we were throwing threads around all the time” (P2). It was easy for this participant to accidentally call into C++ from the wrong Rust thread, but they saw this as less challenging to manage than it would have been to implement the entire application in C++:
@@
Table 3. How often survey respondents would avoid passing abstract data types by value to foreign functions and converting foreign raw pointers into safe references. Percentages on the left include “Always” and “Most of the time”, while those on the right include “Sometimes” and “Never”. The number of respondents in the sample is listed in parentheses on the righthand side.
ID Question Sample Distribution
3a How often do you intentionally avoid passing Rust’s abstract data types (structs, enums) by value across FFI boundaries? Used FFI 49%24%(112)
3b How often do you intentionally avoid converting raw pointers to memory allocated by FFI calls into safe references, such as &T or &mut T? 23%43%(112)
Always    Most of the time    About half the time    Unsure    Sometimes    Never
@@
Participants attempted to keep interoperation with foreign code as straightforward as possible. When declaring bindings, one participant preferred using primitive types and pointers instead of abstract data types like Option and NonNull, even though Rust allows these types to be implicitly cast into raw pointers (Community, 2026b). They felt that these casts hid useful contextual information, so they preferred to use raw pointers explicitly. A few participants indicated that they had avoided passing structs by value across foreign boundaries in certain situations. One of these participants observed that certain Rust programs would behave incorrectly when structs were moved across a foreign boundary instead of copied, or if they had a destructor on the C++ side of the boundary. They typically avoided passing structs by value in either situation, and they recommended that “…if you’re doing FFI, the only things you should pass by value are copy types” (P2). Other interview participants also indicated that they felt comfortable passing structs by value as long as they had copy semantics and equivalent types on each side of the boundary.
@@
We investigated these information-hiding practices using the survey questions shown in Table 0, where respondents indicated if they “intentionally avoid” passing structs by value or casting foreign pointers into Rust’s references. We added “Unsure” as a response option for these questions since they indirectly rely on participants having a notion of why they would “intentionally” avoid these behaviors. Neither practice was consistently followed, but slightly more respondents would avoid passing abstract data types by value most or all of the time (49%, 3), than would avoid casting foreign pointers into references (43%, 3).
Key Findings - RQ1 (Interoperation).
How do Rust developers reason about memory safety across foreign function boundaries?
The majority of participants called foreign functions, which were typically written in C and C++. Certain foreign libraries used aliasing and concurrency patterns that conflicted with the expectations of Rust’s type system, which made them difficult to encapsulate. To make interoperation easier, participants minimized their interaction with foreign libraries, but they frequently used Rust’s memory containers as allocators, and few actively avoided casting foreign pointers into safe references. Developers identified safety properties and design patterns of foreign APIs that match the characteristics of libraries where McCormack et al. (McCormack et al., 2025b) detected aliasing violations, indicating that foreign function calls may be a significant unchecked source of undefined behavior in the Rust ecosystem.
@@ 5.3. RQ2 - Tooling
Table 4. How often survey respondents audited their dependencies, used a debugger, wrote tests, and executed tests in Miri (if they had used Miri at least once). Percentages on the left include “Always” and “Most of the time”, while those on the right include “Sometimes” and “Never”. The number of respondents in each sample is included in parentheses on the righthand side.
@@
4a How often do you write tests for Rust applications that use unsafe? All 61%30%(160)
Used Miri 72%19%(98)
4b How often do you run test cases in Miri? Used Miri 38%56%(98)
4c How often do you audit your dependencies’ use of unsafe? All 6%88%(160)
Used Auditing Tools 10%80%(81)
Always    Most of the time    About half the time    Sometimes    Never
Table 5. Whether or not survey respondents who had used foreign functions would trust the correctness of foreign function bindings written by hand and generated using a tool. Percentages on the left include “Definitely yes” and “Probably yes”, while those on the right include “Probably not” and “Definitely not”. The number of respondents the each sample is included in parentheses on the righthand side.
ID Question Sample Distribution
5a Do you trust FFI bindings that are written by hand? Used FFI 45%7%(112)
5b Do you trust FFI bindings that are generated by a tool? 79%2%(112)
Definitely yes    Probably yes    Might or might not    Probably not    Definitely not
Participants used a wide variety of development tools to assist with validating their design choices. We focused primarily on dynamic bug-finding tools and static tools for generating bindings to foreign functions. However, certain participants also described their experiences using debuggers, verifiers, and auditing tools.
@@
Most participants used dynamic bug finding tools. Interview participants often mentioned Miri (Jung et al., 2026), AddressSanitizer (Serebryany et al., 2012), and Valgrind (Nethercote and Seward, 2007). We prompted survey respondents to indicate which dynamic tools they had used from a list of options mentioned by our interview participants. Survey respondents most frequently used Miri (61%), Valgrind (44%), AddressSanitizer (26%), and cargo-fuzz (Rust Community, 2025a) (26%). Each of the remaining tools listed in our survey was used by less than 15% of respondents. Three interview participants and 31% of survey respondents used some form of fuzzing tools. One interview participant leveraged industry sponsorship to fuzz a JIT compiler for hours-on-end. Another participant implemented a BTree data structure with optimizations for high performance computing, and they used libfuzzer (LLVM Project, 2026) to perform differential testing (McKeeman, 1998) against the implementation provided by Rust’s standard library. Of the 26 developers interviewed by Höltervennhoff et al. (Höltervennhoff et al., 2023) only seven had used Miri, seven used fuzzing tools, and three used Valgrind. Our results indicate that this is not representative of a broader trend in the Rust community.
The majority of survey respondents also wrote tests for applications that used unsafe code. As shown in Table 0, 61% (4) of all participants wrote tests most of the time or always. This was equally true for the subset of participants who had used Miri; 72% would write test cases at least most of the time. However, 56% (4) of survey respondents who had used Miri only sometimes used it to run their test cases, if at all. Höltervennhoff et al. (Höltervennhoff et al., 2023) also had similar findings. The majority of their participants wrote unit tests for unsafe code, but a subset claimed that they would infrequently run these tests because foreign function calls or low-level hardware interaction made testing “impossible.” While notable, we cannot make any causal claims between Miri’s limitations and this lack of testing in our survey population. Additionally, we did not ask if survey respondents regularly ran their test cases. However, interview participants frequently mentioned that Miri’s slow performance and lack of support for foreign function calls prevented them from using it to find bugs in certain categories of applications, and survey respondents who had used Miri encountered similar issues. The majority (62%) were deterred from using it due to one or more of the following issues: lack of support for foreign functions (43%), slow performance (26%), or lack of support for inline assembly (19%). Developers interviewed by Höltervennhoff et al. (Höltervennhoff et al., 2023) had also struggled with Miri’s lack of support for key features, but none had mentioned its performance.
@@
Interview participants encountered several issues with tools that generate bindings to foreign functions. Participants reported that bindgen (Rust Community, 2026a) added significant runtime overhead during build times and was not as expressive as they desired. One participant observed that bindgendid some weird stuff for anonymous unions and structs” (P17) but that this was easy to fix by modifying the tool’s configuration. However, another participant found it necessary to switch from using bindgen to writing bindings by hand, because it was not capable of expanding the macros that they had used to define core data structures. Others felt that the bindings they wrote by hand were more idiomatic than the output of a tool. The majority (58%) of survey respondents who created bindings used multiple methods, while 28% only generated bindings and 14% only wrote them manually.
Interview participants who wrote bindings by hand viewed this process as error-prone, noting that bindings become a “fiddly mess” (P5) when you need to ensure that interfaces remain consistent on each side of the foreign boundary. These errors clearly do occur in practice, as 35% of survey respondents who had called foreign function had also encountered incorrect bindings. McCormack et al. (McCormack et al., 2025b) also found several instances of incorrect bindings in Rust libraries. Some were intentional—developers neglected to add return types for values they did not use—but others introduced aliasing violations or uninitialized reads. Survey respondents who had called foreign functions were somewhat more likely to trust generated bindings. As shown in Table 0, only 45% (5) of respondents would at least “probably” trust handwritten bindings, compared to 79% (5) who would trust generated bindings. However, developers were still more likely to trust handwritten bindings than to distrust them.
@@
Few interview participants reported using a debugger with Rust code. Only 10% of survey respondents used a debugger daily, 29% used one monthly, and 44% used one yearly at most, if at all. One interview participant had used both MSVC’s debugger and LLDB (LLVM Project, 2025), and while both worked, neither matched the quality of the other parts of the Rust toolchain, “where you get all the bells and whistles that you want, and things just work nicely together” (P3). Mozilla’s RR debugger (O’Callahan et al., 2017) was useful for one participant to resolve bugs found from fuzzing, but it was generally more common for participants to report using “printf-style” (P7) debugging.
@@
Few interview participants and only 10 survey respondents used tools that apply formal methods for static and dynamic verification. Interview participants mentioned using Kani (VanHattum et al., 2022), Prusti (Astrauskas et al., 2019), and Crucible (Galois, 2025). Kani was useful for one participant, who needed to prove that a particular routine would never execute more than twice for any input. Another participant who had “tried a bunch of the tools” in this category found that Prusti was the most complete, but they felt that it would be more effective if these tools were “inherently parts of Rust” so that they could “hoist a lot of currently unsafe code into safe code” (P11). Of the ten survey respondents who had used “formal methods tools,” five had used Kani, three had used Prusti, three had used Creusot (Denis et al., 2022), one had used Flux (Lehmann et al., 2023), and one had used Crucible. We provided a short-response option for this question to include any tools that were missing from our list.
@@
Survey respondents rarely audited their dependencies’ use of unsafe code. More than 80% (4) of respondents only sometimes audited their dependencies’ use of unsafe code, if at all. It is possible that a subset of the participants who answered “Never” to this question did not have any dependencies to audit. However, 51% of our respondents had used automated auditing tools, compared to four interviewed by Höltervennhoff et al. (Höltervennhoff et al., 2023). Of this subset, 63% had used cargo-audit (Rust Community, 2025f), 42% had used cargo-update (Rust Community, 2025e), 31% had used cargo-deny (Rust Community, 2025c), 22% had used cargo-geiger (Rust Community, 2025d), and only 7% had used cargo-vet (Rust Community, 2025b).
@@
What tools do Rust developers use when contributing to applications that include unsafe code, and how could tooling be improved?
Miri appears to be the de facto bug finding tool for the Rust ecosystem, since it was used by 61% of all respondents. However, 62% of the respondents who had used Miri before were deterred from using it again due to its performance and lack of support for key features—predominately, foreign function calls. Tools for generating foreign function bindings were often slow and required additional configuration, so a subset of developers still preferred writing bindings by hand. Participants infrequently used debuggers, and only 10 survey respondents had used formal methods tools. Survey respondents infrequently audited their dependencies, but 51% had used one or more auditing tools. Improvements to Miri may be the most effective way to provide stronger safety guarantees for the majority of our audience, since it was the most widely used tool in our survey.